# Structured Probabilistic Inference and Generative Modeling

### v2

@@ Appendix C Ablations
We did several ablations on the quality of the NADEs and their impact on the one-shot causal discovery phase. In particular, Table 2 presents multiple $\text{Tf}_{x},\text{Tf}_{y}$ with respectively 90 and 15 million parameters or 34 and 4 million parameters. We also varied the context window (conditioning set $\boldsymbol{Z}$), trained on different amounts of data (tokens) and reported the classification results on the test set of $\text{Tf}_{y}$ alone. We didn’t output the running time since it was approximately the same for all NADEs: $1.27$ minutes of 50,000 samples and $0.14$ for 5000.
We observe that scaling up the NADEs model size and the trained data show the biggest improvements. After, it is via the context $c$, which after $c=15$ shows declined performance for a larger number of samples. We then choose the backbone with 1.5B Tokens, 105m parameters and a context $c=15$ for our experiments.
@@ C.2 Sampling Number
We experimented with different numbers of $n$ for the sampling method across different averaging (micro, macro, weighted), Fig. 5. We performed 8 different runs and reported the average, standard deviation and elapsed time. We can say that generally, sampling with a bigger $N$ tends to decrease the standard deviation and give more reliable Markov Boundary estimation. Moreover, as we process more samples, the model is gradually improving at a logarithmic growth until it converges to a final score. We also verify that our time complexity is linear with the number of samples $N$. Based on these results, we choose generally $N=68$ as the number of samples.
@@ C.3 Dynamic Thresholding
Figure 6: Evolution of one-shot F1 Score, Precision and Recall in function of coefficient $k$. Results are reported using 1-sigma error bar.
@@ C.5.2 Mutual Information
The BES is the second phase of GES [ges], where edges are removed one after the other to maximise a criterion $S$. Heuristics approaches [approx_fusionPUERTA2021155, delSagrado2001] aim to solve this problem by optimizing:
@@
We would like to reuse the estimated conditional mutual information, Eq. (4) and profit from the parallelised inference of Phase 1 (Fig. 2).
As argued out by JanzingDominik2013QCI, a causal strength measure (or criterion) $C_{X_{i}\rightarrow Y_{j}}$ should possess multiple properties. Notably that if $C_{X_{i}\rightarrow Y_{j}}=0$, then the joint distribution satisfies the Markov condition with respect to the DAG obtained by removing the arrow $X_{i}\rightarrow Y_{j}$}. Moreover, the true DAG reads $X_{i}\rightarrow Y$ iff $C_{X_{i}\rightarrow Y_{j}}=I(X_{i},Y_{j})$.
It comes as a natural criterion to merge edges from multiple causal graphs. However, it remains tricky to estimate [estimating_mi]. Using the chain rule of conditional mutual information [cover1999elements], we can rewrite it as:
@@
Where $I(Y_{j},X_{i},\boldsymbol{Z})$ is the interaction information [cover2006elements], which tells us wether knowing $\boldsymbol{Z}$ explains away the dependency between $X_{i}$ and $Y_{j}$ (negative interaction), or enhances it (positive interaction):
@@
This formulation encourages informative yet parsimonious graph structures, correcting for underrepresented labels via the regularisation term. It is also efficient since CAIG is decomposable [ges] like BIC with the local $s_{I}$. This criterion is denoted as BES mi imbalance par in our experiments in Fig. 4.
@@ Appendix D Implementation
A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], our method is agnostic to both. As illustrated in Figure 2, all steps are parallelised on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels, leading to $\text{BS}\times N\times L$ CI-tests per batch $D=\{S^{0}_{l},\ldots,S^{m}_{l}\}$. Consequently, time complexity transitions from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch due to GPU parallelism. The complexity is bounded by the Transformers’ inference part, where it scales quadratically with the sequence length $\mathcal{O}(L^{2})$ if one uses vanilla self-attention [tf].

### v3

@@ Appendix C Ablations
We did several ablations on the quality of the NADEs and their impact on the one-shot causal discovery phase. In particular, Table 2 presents multiple $\text{Tf}_{x},\text{Tf}_{y}$ with respectively 90 and 15 million parameters or 34 and 4 million parameters. We also varied the context window (conditioning set $\boldsymbol{Z}$), trained on different amounts of data (tokens), and reported the classification results on the test set of $\text{Tf}_{y}$ alone. We didn’t output the running time since it was approximately the same for all NADEs: $1.27$ minutes of 50,000 samples and $0.14$ for 5000.
We observe that scaling up the NADEs model size and the trained data show the most significant improvements. Afterward, it is via the context $c$, which, after $c=15$, shows a decline in performance across a larger number of samples. We then choose the backbone with 1.5B Tokens, 105m parameters, and a context $c=15$ for our experiments.
@@ C.2 Sampling Number
We tested different values of $N$ for the sampling method across averaging methods (micro, macro, weighted), as shown in Fig. 5. We performed eight different runs and reported the average, standard deviation, and elapsed time. In general, sampling with a larger $N$ tends to reduce the standard deviation and yield more reliable Markov Boundary estimates. Moreover, as we process more samples, the model gradually improves, following a logarithmic growth pattern until it converges to a final score. We also verify that our time complexity is linear with the number of samples $N$. Based on these results, we generally select $N=68$ as the sample size.
@@ C.3 Dynamic Thresholding
Figure 6: Evolution of one-shot F1 Score, Precision, and Recall in function of coefficient $k$. Results are reported using 1-sigma error bar.
@@ C.5.2 Mutual Information
The BES is the second phase of GES [ges], where edges are removed one after the other to maximize a criterion $S$. Heuristics approaches [approx_fusionPUERTA2021155, delSagrado2001] aim to solve this problem by optimizing:
@@
We want to reuse the estimated conditional mutual information, Eq. (4), and profit from the parallelized inference of Phase 1 (Fig. 2).
As argued out by JanzingDominik2013QCI, a causal strength measure (or criterion) $C_{X_{i}\rightarrow Y_{j}}$ should possess multiple properties. Notably, if $C_{X_{i}\rightarrow Y_{j}}=0$, then the joint distribution satisfies the Markov condition with respect to the DAG obtained by removing the arrow $X_{i}\rightarrow Y_{j}$}. Moreover, the true DAG reads $X_{i}\rightarrow Y$ iff $C_{X_{i}\rightarrow Y_{j}}=I(X_{i},Y_{j})$.
It is a natural criterion for merging edges across multiple causal graphs. However, it remains tricky to estimate [estimating_mi]. Using the chain rule of conditional mutual information [cover1999elements], we can rewrite it as:
@@
Where $I(Y_{j},X_{i},\boldsymbol{Z})$ is the interaction information [cover2006elements], which tells us whether knowing $\boldsymbol{Z}$ explains away the dependency between $X_{i}$ and $Y_{j}$ (negative interaction), or enhances it (positive interaction):
@@
This formulation encourages informative yet parsimonious graph structures, correcting for underrepresented labels via the regularization term. It is also efficient since CAIG is decomposable [ges] like BIC with the local $s_{I}$. This criterion is denoted as BES mi imbalance par in our experiments in Fig. 4.
@@ Appendix D Implementation
A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], our method is agnostic to both. As illustrated in Figure 2, all steps are parallelized on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels, leading to $\text{BS}\times N\times L$ CI-tests per batch $D=\{S^{0}_{l},\ldots,S^{m}_{l}\}$. Consequently, time complexity transitions from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch due to GPU parallelism. The complexity is bounded by the Transformers’ inference part, where it scales quadratically with the sequence length $\mathcal{O}(L^{2})$ if one uses vanilla self-attention [tf].